OpenStudy (anonymous):

Need help with evaluating this integral: (16-r^2)^1/2r dr dtheta.

OpenStudy (anonymous):

\[\int\limits_{0}^{2\pi}\int\limits_{0}^{2}\sqrt{16-r^2}r dr d \theta\]

OpenStudy (anonymous):

This problem is worked out in my book, theres a certain part where I'm stuck on how they evaluated the problem i'll post that up

OpenStudy (anonymous):

\[-1/3\int\limits_{0}^{2\pi}(16-r^2)^{3/2} \theta \] Where are they getting the squareroot of 3 here? \[-1/3 \int\limits_{0}^{2\pi}(24\sqrt{3}-64)d \theta\]

OpenStudy (anonymous):

I understand the 24 and the 64... oh and I forgot to put the inner limits in brackets at the end of the first equation i typed in.

OpenStudy (anonymous):

firm producing hokey sticks has a production function given by Q=2√(K.L) Where: K= Capital inputs L= Labour inputs In the short run, the firm amounts of capital equipments is Fixed at K=100. The rental rate for capital is V=1.00 Shillings and the wage rate for L=4.00 shillings. Calculate the firms short run costs functions as well as the short run average costs Functions. What is the firms short run marginal costs. any one who can help

OpenStudy (anonymous):

\[\int\limits_{0}^{2\pi}\int\limits_{0}^{2}\sqrt{16-r^2}r dr d \theta\] first \[\int\limits_{0}^{2}\sqrt{16-r^2}r dr\] a u sub gives \[-\frac{1}{3}(16-r^2)^{\frac{3}{2}}\]

OpenStudy (anonymous):

yep, I follow you so far

OpenStudy (anonymous):

replace x by 2 and get \[-\frac{1}{3}(16-2^2)^{\frac{3}{2}}\] \[-\frac{1}{3}\sqrt{12}^3\] \[-\frac{1}{3}\times 24\sqrt{3}\] \[-8\sqrt{3}\]

OpenStudy (anonymous):

replace x by 0 and get \[-\frac{1}{3}(16)^{\frac{3}{2}}\] etc

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

wait, now I'm confused about the 24. Where is that coming from?

OpenStudy (anonymous):

Do you still don't know where the 24√3 comes from?

OpenStudy (anonymous):

Yeah

OpenStudy (dumbcow):

\[\sqrt{12}^{3} = (2\sqrt{3})^{3} = 8*3*\sqrt{3} = 24\sqrt{3}\]

OpenStudy (anonymous):

That other three is coming from -1/3 right?

OpenStudy (anonymous):

if so then i think i get it

OpenStudy (anonymous):

No i'm wrong, cause then it would have been -24squareroot of 3

OpenStudy (dumbcow):

? not following you

OpenStudy (anonymous):

I'm a bit confused on your answer, I know you got 8 from 2^3 but how are you getting the 8*3 part?

OpenStudy (dumbcow):

you should understand how they get to this point now \[-1/3\int\limits_{0}^{2\pi}(24\sqrt{3}-64) d \theta\]

OpenStudy (dumbcow):

\[\sqrt{3}^{3} = \sqrt{3}*\sqrt{3}*\sqrt{3} = \sqrt{9}*\sqrt{3}=3*\sqrt{3}\]

OpenStudy (anonymous):

oh ok now I see. Thanks.

OpenStudy (dumbcow):

no problem